# LUMA: A Benchmark Dataset for Learning from Uncertain and Multimodal Data

### v2

@@ 3.3 Text Modality
For text modality, the main constraint was that the text segments had to talk about the subject of the images. For that, we decided to employ a generative model, and generate text segments about the class label. We utilized Google’s Gemma-7B Instruct model (Mesnard et al., 2024) to generate more than 1,200 texts samples per class, using 13 different prompts. Here is an example of a prompts used for generation (please find the full list of the prompts in the supplementary material).
"You are explaining a five year old child what the word <word> means. Use very simple and explanatory language, so the kid will understand the meaning of the word <word>. Tone: Casual, complexity: simple"
Gemma-7B Instruct was chosen, since according to their technical report (Mesnard et al., 2024), it outperforms other open LLMs with similar size, in 11 out of 18 tasks. Moreover, in our experiments, it provided better answers to our prompts compared with Mistral-7B (Jiang et al., 2023).
Figure 3: Text generation and validation pipeline
To validate that the generated texts accurately represent the labels, we masked all label occurrences in the text and fed the masked text back into the Gemma-7B Instruct model, asking it to classify the text into one of the labels (see the prompt in the supplementary material). Based on the prediction of the model, if the prediction matched the ground truth label, we accepted the sample as validated. In total, we accepted 55,953 text samples.
After manually analyzing some of the generated texts, we noticed that there were samples with offensive biases and stereotypes (Some examples are included in the supplementary material). Particularly, we noticed a lots of gender bias for classes "man", "woman", "boy" and "girl". To find the proportion of the biased data, we asked the Gemma model to find out if the given text contains gender, racial, religious, or cultural biases. We found out, that indeed, the aforementioned 4 classes have a huge amount of gender bias (see the bias detection statistics in the supplementary material). Our hypothesis is, that describing a man or woman in an unbiased way is a challenging task for LLM models (as well as for humans), which are trained on unbalanced data (Kotek et al., 2023). For that purpose, we reconstruct the prompts (see in the supplementary material), to explicitly provide topics and keywords with occupations, which will minimize the bias. We then rerun the bias detection prompt, which found fewer biased samples and allowed us to filter enough samples to be included in the final dataset.
3.4 Dataset compilation
Based on the collected samples from the 3 modalities (image, text, audio), we wanted to compile a dataset with little uncertainties, and later, provide tool to inject uncertainties on demand. Our priority was to propose several options for uncertainty control and parameter setting: data diversity, sample noise, label noise, OOD injection.
3.4.1 Data Diversity
Figure 4: t-SNE (van der Maaten & Hinton, 2008) visualization of audio data points from class "man" (in gray), and sampled points with different diversity parameter $k$. The higher is the value of $k$, the more concentrated (less diverse) the points are.
With a fixed number of data points, increasing the diversity of the data enhances the information passed to the model, thereby it shall reduce the epistemic uncertainty. Conversely, when samples are concentrated at a single point in the latent space, they encode less information, which shall lead to greater epistemic uncertainty in areas where data is scarce (Figure 5). Hence, controlling the diversity of the data allows us to study the behavior of epistemic uncertainties under varying amounts of information.
Figure 5: t-SNE (van der Maaten & Hinton, 2008) visualization of audio data points for all classes, sampled with different diversity parameter $k$. With higher $k$ we have more concentrated samples, and more separation between classes. The diversity can similarly be controlled for the other modalities.
To control the diversity, we extract deep features from each modality (Wav2Vec (Baevski et al., 2020) for audio, BERT (Devlin et al., 2019b) for text and VGG-11 (Simonyan & Zisserman, 2015) for images), and compute the inverse distance of each sample to the center (mean vector) of its class, raised to the power of $k$:
$D_{i}=\frac{1}{\left\|F_{i}-\frac{1}{|C|}\sum_{j\in C}F_{j}\right\|_{2}^{k}},\ \ i\in C,$ (1)
where $F$ represents the deep feature vectors extracted from the samples, $C$ is the set of data indices belonging to the class, and ||$\cdot$|| ⋅ | measures the cardinality of the set. Then, having the inverse distances, we sample points from categorical distribution $x_{n}\sim Categorical(D)$. In Eq. 1, $k$ is the variable controlling the diversity. If $k=0$ the sampling is uniform. The bigger $k$, the higher probability of selection will be applied to the samples closer to the center.
Sample Noise: We want to have an option to inject controlled amount of noise in the data. This may reduce the information in each data sample, and increase the classification difficulty. With our hypothesis, this may affect both epistemic and aleatoric uncertainty degrees. This type of noise can also be very beneficial for estimating the model robustness to noise. We apply different types of noise to each modality.
For audio modality, we added background noise from ESC-50 dataset (Piczak, 2015) to each sample, using the audiomentations [^10] https://github.com/iver56/audiomentations library. The amount of the minimum and maximum signal-to-noise ratio, as well as the proportion of the noisy data is set as a hyper-parameter.
For text modality, we utilize the nlpaug (Ma, 2019) library, to add different types of noise. The user has option to choose a subset of noise types from: 1) Keyboard noise that simulates keyboard distance error; 2) OCR noise that simulates OCR engine noise; 3) Random character noise to insert, substitute, or delete random characters; 4) Antonym noise to swap random words with their antonyms; 5) Random word noise to insert, substitute, or delete random words; 6) Spelling noise to add spelling mistakes according to spelling mistake dictionary; and 7) Back-translation noise to translate the text to another language, and then translate back to English. The parameters of these noise types can be specified by the user, and are transferred to the nlpaug library for adding the specific type of noise.
For image modality, we added different types of noise suggested and implemented by Hendrycks & Dietterich (2018). 15 perturbations are included such as: adding Gaussian noise, shot noise, impulse noise, defocus blur, frosted glass blur, motion blur, zoom blur, snow, frost, fog, changing the brightness, contrast, elasticity, pixelating, and JPEG compressing.
Label Noise: Another way to insert aleatoric uncertainty is to add label noise in the data (i.e., randomly switch the labels of some samples). Since the uncertainty induced by this type of noise cannot be reduced with additional amount of data, this noise shall increase the aleatoric uncertainty of the model. To insert this type of noise, we choose samples at random (according to a probability defined by the user), then, based on their deep features (Wav2Vec (Baevski et al., 2020) for audio, BERT (Devlin et al., 2019b) for text and VGG-11 (Simonyan & Zisserman, 2015) for images), we find the average distance to the 5 nearest points from each class, and assign the label of the class that has the closest mean distance.
OOD Injection: Ideally, the models shall be uncertain on data points from unknown distribution (i.e., distribution they haven’t been trained on). In the literature, often the OOD samples are taken from another dataset, which can simplify the problem, because such samples are far from the training data. For this matter, we kept a separate set of samples from the same dataset, but belonging to classes that are not present in the training data, as OOD samples.
4 Baseline models with Uncertainty Quantification
We develop a baseline model with three different uncertainty quantification algorithms, to serve as a starting point for other research and benchmarking initiatives. For the sake of simplicity, we choose late or decision fusion approaches, where we have classification networks for each modality, and then fuse their decision by simple averaging the output logits. The implementation details of the architectures can be found in the supplementary material.
For uncertainty quantification, we implemented 3 approaches: Monte Carlo Dropout (MCD) (Gal & Ghahramani, 2016), Deep Ensemble (DE) (Lakshminarayanan et al., 2017), Reliable Conflictive Multi-View Learning (RCML) (Xu et al., 2024).
In Monte Carlo Dropout and Deep Ensembles, we use the aleatoric entropy and the epistemic entropy as as uncertainties measures $H_{\text{Ale }}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{\text{Ale }}(y\mid\mathbf{x})\right)$ and $H_{\mathrm{Epi}}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{Epi}(y\mid\mathbf{x})\right)$, where $p_{Epi}$ and $p_{Ale}$ are the probabilities obtained according to (Valdenegro-Toro & Mori, 2022). In RCML, we measure the aleatoric uncertainty with the expected entropy such as:
$\mathbb{E}_{p(\bm{\pi}\mid\mathbf{x},\hat{\bm{\theta}})}[H[P(y\mid\bm{\pi})]]=-\sum_{k=1}^{K}\frac{\alpha_{k}}{\alpha_{0}}\left(\psi\left(\alpha_{k}+1\right)-\psi\left(\alpha_{0}+1\right)\right),$ (2)
where $\alpha_{k}$ is the $k$-th concentration parameter of the Dirichlet distribution, and $\alpha_{0}$ is the sum of all concentration parameters. $\psi$ is the digamma function. As a measure for epistemic uncertainty, we take $\frac{N}{\alpha_{0}}$, where $N$ is the number of classes.
We evaluate the measures of accuracy and uncertainty of the models on the clean dataset, the dataset with reduced diversity ($\searrow$ Diversity), the dataset with increased sample noise ($\searrow$ Sample Noise), and the dataset with switched label $\nearrow$ Label Noise). The results are summarized in Tables 1. (For noise generation parameters, please refer to the supplementary material). Since in RCML the uncertainty is quantified differently compared with MCD and DE, we cannot directly compare their values, and relative changes to the two type of uncertainty compared to the clean dataset is reported in the table. We observe that accuracy always decreases with increasing the label noise, but reducing diversity and increasing sample noise may not always decrease accuracy in the image modality.
Table 1: Results for UQ with baseline models. The absolute values are reported for clean dataset, and changes in percentages relative to clean dataset are reported for the noisy versions of LUMA dataset.
Method Clean $\searrow$ Diversity $\nearrow$ Label Noise $\nearrow$ Sample Noise
Ale. Epi. Ale. Epi. Ale. Epi. Ale. Epi.
MCD Image 1.00 1.03 -15.73% -11.66% +59.20% +54.51% +4.44% +2.18%
MCD Audio 0.52 0.70 -5.54% +2.16% +96.63% +54.49% +23.12% +14.40%
MCD Text 0.37 1.01 -3.91% -2.62% +93.59% +2.41% +64.96% -2.03%
MCD Multi. 0.26 0.78 -8.52% -1.21% +122.44% +11.60% +59.14% +9.89%
DE Image 1.45 1.40 -37.49% -8.54% -7.43% +0.24% -18.46% -3.22%
DE Audio 0.56 0.99 -27.39% -3.34% +156.40% +50.43% +70.26% +34.41%
DE Text 0.42 1.01 +5.02% -6.15% +81.26% -0.51% +62.24% -7.11%
DE Multi. 0.31 0.82 -22.80% -3.40% +115.15% +20.62% +45.97% +5.54%
Table 2: OOD Detection AUC Values for Different Methods
Method MCD DE RCML Multi.
Image Audio Text Multi. Image Audio Text Multi.
AUC 0.54 0.47 0.53 0.50 0.54 0.49 0.54 0.50 0.91
Figure 6: Changes in uncertainty estimations under different proportions of label noise (in percentages). The RCML (Reliable Conflictive Multi-View Averaging) approach consistently increases both aleatoric and epistemic uncertainties with increased label noise. In contrast, the MCD (Monte Carlo Dropout) and DE (Deep Ensemble) models sometimes fail to increase the corresponding uncertainty estimations in this experiment.
We evaluate AUC score for OOD detection by the networks based on the epistemic uncertainty. The results are summarized in Table 2. We can see that Monte Carlo Dropout and Deep Ensembles fail to provide epistemic uncertainty values suitable for OOD detection in LUMA dataset, with a poor performance of approximately 0.5 AUC value. On the other hand, the RCML achieves an outstanding AUC score of 0.91, indicating that the epistemic uncertainty values quantified with this method can be effectively used for OOD detection.
To further evaluate the qualities of the uncertainties of the different models, we estimate the epistemic and aleatoric uncertainties under different amounts of label noise. Ideally, we expect a good uncertainty quantification algorithm to provide higher uncertainty values for more noisy data. As we can see from Figure 6, only RCML consitently raises the uncertainty estimates under increased label noise, which again shows the higher quality of its uncertainty estimates over the other baselines.
5 Conclusion
In this paper, we propose LUMA, a new multimodal dataset for learning from uncertain and multimodal data and benchmarking. It includes image, audio and text modalities and a Python package for compiling different versions of the dataset with various amounts and types of uncertainty and noise.
The dataset can be easily extended with additional modalities and augmented with more data samples. The open-source nature of the data compilation pipeline and code for uncertainty and noise generation facilitates the integration of new contributions from the community to promote multimodal uncertainty studies and benchmarking initiatives.
Appendices
Appendix A Dataset links and Licenses
In this section we present the dataset links and licenses for the LUMA dataset.
Appendix B Details about Data Collection
B.1 Image Modality
B.1.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we use the CIFAR-100 and CIFAR-10 (Krizhevsky, ) datasets and EDM (Karras et al., 2022) generated images from the same distribution for the image modality.
Unfortunately, we failed to find an explicit license for the CIFAR-10/100 datasets, hence, we did not include the images in our dataset repository, and instead, allow the users to download them directly using the provided LUMA dataset compilation and uncertainty generation tool.
The EDM Generated images retrieved from https://huggingface.co/datasets/P2333/DM-Improves-AT are published under the Apache-2.0 license.
B.2 Audio Modality
In this section, we will present some additional information regarding data collection process of audio modality.
B.2.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we collected audio data from 3 sources: The Spoken Wikipedia (Köhn et al., 2016), LibriSpeech (Panayotov et al., 2015), and Mozilla Common Voice (Ardila et al., 2020) corpora.
The Spoken Wikipedia dataset is published under the CC BY-SA 4.0 license. Following the requirements of this license, we also distribute our dataset under the same license. The audio samples extracted from this dataset are located in the sw_audio directory in LUMA dataset.
The LibriSpeech dataset is published under the CC BY 4.0 license. The audio samples extracted from this dataset are located in the ls_audio directory in LUMA dataset.
The Mozilla Common Voice dataset is published under the CC0 (public domain) license. The audio samples extracted from this dataset are located in the cv_audio directory in LUMA dataset.
B.2.2 Audio extraction and validation
To extract the audio pronunciation of each word as described in Section 3.2, we follow the Algorithm 1. First, the algorithm iterates through all the words in the transcripts. If a word matches one of the labels or their plural forms, it is added to the extraction_words list (Lines 3-10), along with its corresponding time steps and audio path. Next, the algorithm processes each word in this list, using the extract_audio function to extract the relevant audio segments. These segments are then added to the extracted_audios list, which is returned as output of the algorithm (Lines 11-17).
Algorithm 1 Audio Extraction Algorithm
1:  Input: labels $\leftarrow$ list of the CIFAR-100/10 labels
2:  Input: transcripts_paths $\leftarrow$ list of (transcript, audio_path) tuples
3:  extraction_words $\leftarrow$ empty list
4:  for transcript, audio_path in transcripts_paths do
5:     for word in transcript do
6:        if word in labels or word + ’s’ in labels or word + ’es’ in labels then
7:           extraction_words.append((word, transcript[word].start_time, transcript[word].end_time, audio_path))
8:        end if
9:     end for
10:  end for
11:  extracted_audios $\leftarrow$ list
12:  for entry in extraction_words do
13:     word, start_time, end_time, path $\leftarrow$ entry
14:     audio_segment $\leftarrow$ extract_audio(path, start_time, end_time)
15:     extracted_audios.append(audio_segment)
16:  end for
17:  return  extracted_audios
Then, we perform automatic validation with the OpenAI’s Whisper Large V3 model (Radford et al., 2023), and on the subset of the negative predictions of the model we perform a manual validation. For the manual validation, we have collected manual annotations thanks to the efforts of our volunteering colleagues and friends. To build the annotation interface, we opted for the Label Studio (Tkachenko et al., 2020-2022), an open-source data labeling platform. You can see the screenshot of the annotation interface in Figure 7. In total, we collected 16,744 manual annotations from 17 human annotators over a period of 2 months. Each sample was annotated by 2 annotators, and only annotations with 100% agreement were accepted.
After automatic and manual validation, we have 135,096 audio samples with the class distribution shown in Figure 8.
Figure 7: The labeling interface using Label Studio
Figure 8: Number of validated audio samples per class. We will include the classes with higher than 600 samples as in-distribution data, and others as out-of-distribution data.
B.3 Text Modality
For the text modality, as described in the Section 3.3, we employed Google’s Gemma-7B Instruct model (Mesnard et al., 2024) to generate more than 1,200 text samples per class, using 13 different prompts. The 13 prompts are as follows:
@@ B.3 Text Modality
The <word> was replaced with the class labels. After generating the text data, to check if the generated text has any bias or stereotype, we again used the Gemma-7B Instruct model with the following prompt:
Your job is to identify biases in texts. You will be given a text, and you need to classify it into one of: [gender_bias, cultural_bias, racial_bias, religous_bias, no_bias].
@@
The <text> was replaced with the text that needed to be checked. The number of biased and unbiased texts in each class according to Gemma model can be found in Figure 9. As we can see, the labels man, woman, boy, and girl have high amount of biased texts. Some examples of identified biased texts are:
@@
Man, a force of might,
A guardian, protector, and light.
With strength and wisdom, they stand tall,
Strong and proud, answering the call.
@@
Figure 9: The amount of texts with different biases according to Gemma-7B Instruct model.
Figure 10: The amount of texts with different biases according to Gemma-7B Instruct model, after reconstructing the prompts for these 4 labels. Although there is still high amount of bias, we can filter them out and still have enough unbiased texts (more than 800 text passages per class).
To alleviate this issue, for the labels of man, woman, boy and girl, we reconstructed the prompts for text generation. The prompt used is the following:
@@
To validate that the generated text passages could be correctly classified, we fed them back into the Gemma model using the following prompt:
In the text: "<text>", you need to understand what is the <masked_word> or find out what is the whole text about.
You need to classify the text into one of these classes: "<classes>"
@@
Appendix C Details about Baseline architectures and Datasets

### v3

@@ 3.3. Text Modality
For text modality, the main constraint was that the text segments had to talk about the label of the images. For that, we decided to employ a generative model, and generate text segments about the class label. We utilized Google’s Gemma-7B Instruct model (Mesnard et al., 2024) to generate more than 1,200 texts samples per class, using 13 different prompts. The 13 prompts are as follows:
@@
The <word> was replaced with the class labels. Gemma-7B Instruct was chosen, since according to their technical report (Mesnard et al., 2024), it outperforms other open LLMs with similar size, in 11 out of 18 tasks. Moreover, in our experiments, it provided responses that we found more useful for our intended applications than those from Mistral-7B (Jiang et al., 2023).
Figure 4. Text generation and validation pipeline
@@
Figure 4: Fully described in the text.
Figure 5. The amount of texts with different biases according to Gemma-7B Instruct model.
[…]
Woman, Man, Girl and Boy classes have the highest amount of biased texts, while other classes have very low amount of biases.
To validate that the generated texts accurately represent the labels, we masked all label occurrences in the text and fed the masked text back into the Gemma-7B Instruct model, asking it to classify the text into one of the labels. Based on the prediction of the model, if the prediction matched the ground truth label, we accepted the sample as validated. In total, we accepted 55,953 text samples.
After manually analyzing some of the generated texts, we noticed that there were samples with offensive biases and stereotypes, such as:
@@
Particularly, we noticed a lots of gender bias for classes "man", "woman", "boy" and "girl". To find the proportion of the biased data, we asked the Gemma model to find out if the given text contains gender, racial, religious, or cultural biases. We found out, that indeed, the aforementioned 4 classes have a huge amount of gender bias (See Figure 5). Our hypothesis is, that describing a man or woman in an unbiased way is a challenging task for LLM models (as well as for humans), which are trained on unbalanced data (Kotek et al., 2023). Although the model identified a high level of bias in these classes, indicating a potential for LLM-based bias detection, we are unable to assess how accurate or consistent these detections truly are.
Figure 6. The amount of texts with different biases according to Gemma-7B Instruct model, after reconstructing the prompts for these 4 labels. Although there is still high amount of bias, we can filter them out and still have enough unbiased texts (more than 800 text passages per class).
[…]
Regenerated texts have enough unbiased samples to be included in the dataset.
To reduce the biases for the labels of man, woman, boy and girl, we reconstructed the prompts to explicitly provide topics and keywords with occupations, which will minimize the bias. The prompt used is the following:
@@
3.4. Dataset compilation
Based on the collected samples from the 3 modalities (image, text, audio), we first compiled a clean version of the dataset with minimal uncertainty. Our goal was to then provide tools that allow uncertainty to be introduced on demand. We prioritized flexibility, offering multiple options for controlling uncertainty through adjustable parameters such as data diversity, sample-level noise, label corruption, and the injection of out-of-distribution (OOD) samples.
Figure 7. t-SNE (van der Maaten and Hinton, 2008) visualization of audio data points from class "man" (in gray), and sampled points with different diversity parameter $k$. The higher is the value of $k$, the more concentrated (less diverse) the points are.
@@ 3.4. Dataset compilation
Figure 7: Fully described in the caption and text.
3.4.1. Data Diversity
With a fixed number of data points, increasing the diversity of data enhances the information passed to the model; thereby, it shall reduce the epistemic uncertainty. Conversely, when samples are concentrated at a single point in the latent space, they encode less information, which shall lead to greater epistemic uncertainty in areas where data is scarce (Figure 8). Hence, controlling the diversity of the data allows us to study the behavior of epistemic uncertainties under varying amounts of information.
Figure 8. t-SNE (van der Maaten and Hinton, 2008) visualization of audio data points for all classes, sampled with different diversity parameter $k$. With higher $k$ we have more concentrated samples, and more separation between classes. The diversity can similarly be controlled for the other modalities.
@@ 3.4.1. Data Diversity
Figure 8: Fully described in the caption and text.
To control the diversity, we extract deep features from each modality (Wav2Vec (Baevski et al., 2020) for audio, BERT (Devlin et al., 2019a) for text and VGG-11 (Simonyan and Zisserman, 2015) for images), and compute the inverse distance of each sample to the center (mean vector) of its class, raised to the power of $k$:
(1) $D_{i}=\frac{1}{\left\|F_{i}-\frac{1}{|C|}\sum_{j\in C}F_{j}\right\|_{2}^{k}},\ \ i\in C,$
where $F$ represents the deep feature vectors extracted from the samples, $C$ is the set of data indices belonging to the class, and $|\cdot|$ measures the cardinality of the set. Then, having the inverse distances, we sample points from the categorical distribution $x_{n}\sim\text{Cat}(D)$. In Eq. 1, $k$ is the variable controlling the diversity. If $k=0$ the sampling is uniform. The bigger $k$, the higher probability of selection will be applied to the samples closer to the center.
3.4.2. Sample Noise
We want to have an option to inject a controlled amount of noise into the data. This may reduce the information in each data sample and increase the classification difficulty. With our hypothesis, this may affect the aleatoric uncertainty degrees. This type of noise can also be very beneficial for estimating the model’s robustness to noise. We apply different types of noise to each modality.
For audio modality, we added background noise (Human non-speech sounds, urban noises, animal sounds, natural soundscapes, etc.) from the ESC-50 dataset (Piczak, 2015) to each sample, using the audio- mentations [^11] https://github.com/iver56/audiomentations library. The amount of the minimum and maximum signal-to-noise ratio, as well as the proportion of the noisy data, is set as a hyper-parameter.
For text modality, we utilize the nlpaug (Ma, 2019) library, to add different types of noise. The user has the option to choose a subset of noise types from: 1) Keyboard noise that simulates keyboard distance error; 2) OCR noise that simulates OCR engine noise; 3) Random character noise to insert, substitute, or delete random characters; 4) Antonym noise to swap random words with their antonyms; 5) Random word noise to insert, substitute, or delete random words; 6) Spelling noise to add spelling mistakes according to the spellingling mistake dictionary; and 7) Back-translation noise to translate the text to another language, and then translate back to English. The parameters of these noise types can be specified by the user, and are transferred to the nlpaug library for adding the specific type of noise.
For image modality, we added different types of noise suggested and implemented by Hendrycks and Dietterich (2018). 15 perturbations are included such as: adding Gaussian noise, shot noise, impulse noise, defocus blur, frosted glass blur, motion blur, zoom blur, snow, frost, fog, changing the brightness, contrast, elasticity, pixelating, and JPEG compressing. Additionally, common transformations such as cropping or skewing are not part of the default set but can be easily applied by the user outside the framework if needed.
3.4.3. Label Noise
Aleatoric uncertainty can also be introduced by injecting label noise into the dataset, i.e., by randomly altering the labels of certain samples. Since this type of uncertainty arises from inherent noise in the data itself, it cannot be mitigated by increasing the dataset size and thus directly contributes to the model’s aleatoric uncertainty.
To inject label noise in a more structured manner, we randomly select a subset of samples based on a user-defined probability. For each selected sample, we compute its deep feature representation using modality-specific encoders: Wav2Vec (Baevski et al., 2020) for audio, BERT (Devlin et al., 2019a) for text, and VGG-11 (Simonyan and Zisserman, 2015) for images. We then measure the sample’s average distance to the five nearest samples from each class. The sample is reassigned the label of the class with the lowest mean distance, ensuring the new label is still semantically or perceptually close to the original, thereby creating realistic label noise.
3.4.4. OOD Injection
Ideally, the models shall be uncertain on data points from unknown distribution (i.e., distribution they haven’t been trained on). In the literature, often the OOD samples are taken from another dataset, which can simplify the problem, because such samples are far from the training data. For this matter, we kept a separate set of samples from the same dataset, but belonging to classes that are not present in the training data, as OOD samples.
4. Baseline models with Uncertainty Quantification
Table 1. Results for UQ with baseline models. The absolute values are reported for clean dataset, and changes in percentages relative to clean dataset are reported for the noisy versions of LUMA dataset.